# supplementary

### v1

@@ 5 Conclusion
    Appendix A — Proof of Lemma
    • A.1 — Proofs
    • A.2 — Technical Lemma 2
    • A.3 — Proof of Theorem 1
    • A.4 — Convergence Analysis for Label Smoothing
@@
    Appendix B — Implementation Details
@@
    Appendix C — Experimental Details
    • C.1 — Additional Results under the Cross-Dataset Generalization Setting
    • C.2 — Additional Results under the Domain Generalization Setting
    • C.3 — Full Results of the Base-to-New Generalization Setting
    • C.4 — Results under the Few-shot Classification Setting
  4. 4.
    Appendix D — Additional Experiments
    • D.1 — Sensitivity Analysis
    • D.2 — Analysis of the Soft Labels Generated by CSL and ISL
    • D.3 — Combination of CSI and ISL
    • D.4 — Further Analysis of the Alternating Period $K$
    • D.5 — Computational Cost
  5. 5.
    Appendix E — Limitations
Appendix A Proofs
@@
A.1 Proof of Lemma 1
Recall Lemma 1 in Section 3:
Lemma 1.
@@ A.2 Technical Lemma 2
A.3 Proof of Theorem 1
Here we first restate Theorem 1 in Section 3 as follows:
Theorem 1.
@@ A.3 Proof of Theorem 1
Proof. Recall Assumption 1 says the objective function $F(\mathbf{v})$ is $\beta$-smooth, i.e., $\left|\nabla F(\mathbf{v})-\nabla F(\mathbf{u})\left|\leq\beta\right|\mathbf{v}-\mathbf{u}\right|$, which has an equivalent expression (Nesterov 2013):
@@
A.4 Convergence Analysis for Label Smoothing
In this section, we include the convergence analysis for label smoothing derived by (Xu et al. 2020). Note here the context vector $\mathbf{v}_{t}$ are the only learnable parameters, slightly different from the version in (Xu et al. 2020) where all the parameters $\mathbf{x}_{t}$ in the model are learnable.
Lemma 3.
@@ Appendix B Implementation Details
The textual prompts for different datasets used in CSL and ISL follow (Radford et al. 2021; Zhou et al. 2022b), as shown below:
@@
Appendix C Experimental Details
We adopt the same data split strategy as CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a) across all datasets and settings.
Due to the page limit in the main body, we provide detailed results in this section, including additional baselines (i.e., CoCoOp (Zhou et al. 2022a), DePT (Zhang et al. 2024), IVLP (Rasheed et al. 2023), and MaPLe (Khattak et al. 2023a)) integrated with the proposed ATLaS method under the cross-dataset generalization and domain generalization settings, and the full results under both the base-to-new generalization and few-shot classification settings.
Source Target
Table A5: Performance of various methods under the cross-dataset generalization setting. The number in the parentheses denotes the improvement over the baseline method.
C.1 Additional Results under the Cross-Dataset Generalization Setting
In this section, we validate the cross-dataset generalization capability of the proposed ATLaS method by integrating several baseline methods with ATLaS and its two variants, including ATLaS-CSL and ATLaS-ISL under the cross-dataset generalization setting. Results in Table A5 show that integrating ATLaS and its variants consistently improve the cross-dataset generalization performance over baseline methods. Notably, the effectiveness of different variants of ATLaS varies across different baselines. For instance, CoOp and MaPLe exhibit the best results with ATLaS-CSL, CoCoOp and DePT benefit most from ATLaS-ISL, while IVLP shows the best results with the standard ATLaS method.
C.2 Additional Results under the Domain Generalization Setting
Source Target
ImageNet -V2 -Sketch -A -R Avg.
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
+ ATLaS 71.79 64.81 48.76 50.66 76.99 60.31 (+1.03)
+ ATLaS-CSL 71.43 64.27 49.02 51.24 76.85 60.35 (+1.07)
+ ATLaS-ISL 71.72 64.85 48.70 50.50 76.21 60.07 (+0.79)
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.90
+ ATLaS 70.99 64.21 49.13 51.23 76.99 60.39 (+0.49)
+ ATLaS-CSL 71.05 64.38 49.19 51.41 77.08 60.52 (+0.62)
+ ATLaS-ISL 70.80 64.10 48.90 51.30 77.00 60.33 (+0.43)
DePT 72.79 63.94 45.79 46.51 73.34 57.40
+ ATLaS 72.87 64.03 46.90 47.59 74.48 58.25 (+0.86)
+ ATLaS-CSL 72.83 64.11 46.81 47.63 74.19 58.19 (+0.79)
+ ATLaS-ISL 72.83 64.09 46.85 47.57 74.12 58.16 (+0.76)
IVLP 72.40 65.20 48.80 48.00 76.50 59.63
+ ATLaS 72.60 65.20 49.20 48.70 76.70 59.95 (+0.33)
+ ATLaS-CSL 72.40 65.10 49.20 49.00 76.40 59.93 (+0.30)
+ ATLaS-ISL 71.90 65.30 49.50 48.80 76.60 60.05 (+0.42)
MaPLe 72.01 64.70 48.66 49.21 76.31 59.72
+ ATLaS 71.50 64.53 48.87 50.17 76.97 60.14 (+0.42)
+ ATLaS-CSL 71.47 64.53 48.83 50.23 76.70 60.07 (+0.35)
+ ATLaS-ISL 71.30 64.53 49.30 50.20 77.03 60.27 (+0.55)
Table A6: Performance of various methods under the domain generalization setting. The number in the parentheses denotes the improvement over the baseline method.
In this section, we validate the domain generalization capability of the proposed ATLaS method by integrating additional baseline methods with ATLaS and its two variants under the domain generalization setting. Results in Table A6 show that the prompts learned by the proposed ATLaS method and its variants are more domain-generalizable. Those results substantiate the flexibility and effectiveness of the proposed ATLaS method, which can seamlessly integrate with existing prompt tuning methods, enhancing prompt generalization.
C.3 Full Results of the Base-to-New Generalization Setting
In this section, we provide the base-to-new generalization performance on 11 datasets and additionally integrate our ATLaS and its variants with the state-of-the-art PromptSRC method (Khattak et al. 2023b). The results in Tables A7, A8 and A9 show that the proposed ATLaS method can consistently improve the new-class accuracy and harmonic mean accuracy over baseline methods, highlighting its effectiveness in enhancing prompt generalization. However, we observe limited performance gains when applying ATLaS to PromptSRC, potentially attributable to both methods leveraging CLIP’s multi-modal capabilities to capture and exploit inter-class and instance-class relationships. Nonetheless, the proposed ATLaS method achieves performance improvements for PromptSRC on all the datasets.
@@ C.3 Full Results of the Base-to-New Generalization Setting
C.4 Results under the Few-shot Classification Setting
In the few-shot classification setting, prompts are trained using a limited number of labeled samples and evaluated on the remaining samples. We conduct few-shot classification using 1, 2, 4, 8, and 16 samples per class. The indicator function $\delta$ is always set to 1 under the few-shot classification setting. The performance of CoOp and its enhanced variants (e.g., CoOp with ATLaS, ATLaS-CSL, and ATLaS-ISL) on the 11 datasets is shown in Figure A4. It can be observed that CoOp with ATLaS, ATLaS-CSL, and ATLaS-ISL all outperform CoOp in terms of the average performance on the 11 datasets. Those results demonstrate the strong generalization capability of the proposed ATLaS method when the samples of downstream tasks are limited.
Figure A4: Performance of various methods on the 11 datasets under the few-shot classification setting.
Appendix D Additional Experiments
D.1 Sensitivity Analysis
In this section, we investigate the impact of four hyper-parameters (i.e., $\theta$, $\tau_{c}$, $\alpha$, and $K$) on the average performance across 11 datasets under the base-to-new generalization setting and the results are shown in Figure A5.
Figure A5: The average performance of ATLaS (a), ATLaS-CSL (b), ATLaS-ISL (c, d) on all datasets with respect to hyper-parameters (i.e., $\theta$, $\tau_{c}$, $\alpha$, and $K$) under the base-to-new generalization setting.
Impact of the smoothing parameter $\theta$. The smoothing parameter $\theta$ in Eq. (3) balances one-hot labels against the uniform distribution. Figure 5(a) shows the average performance of CoOp with ATLaS across the 11 datasets for $\theta$ values in {0.1,…,0.6}. As $\theta$ increases, soft labels exhibit reduced confidence in target classes and increased attention to non-target classes. This leads to a gradual decline in the base-class accuracy and improvement in the new-class accuracy, while the harmonic mean accuracy remains relatively stable.
Impact of the temperature parameter $\tau_{c}$. The temperature parameter $\tau_{c}$ in Eq. (4) controls the smoothness of the similarity distribution across inter-class textual embeddings. A lower $\tau_{c}$ sharpens the distribution, enhancing sensitivity to inter-class distinctions, while a higher value smooths the distribution. Figure 5(b) shows the average performance of CoOp with ATLaS-CSL across 11 datasets for $1/\tau_{c}$ values in {10, 20,…, 100}. It can be seen that as $\tau_{c}$ decreases, the base-class accuracy initially increases and subsequently stabilizes, while both the new-class accuracy and the harmonic mean accuracy exhibit an initial increase followed by a decline.
Impact of the rectification coefficient $\alpha$. The rectification coefficient $\alpha$ in Eq. (5) controls how significantly the one-hot label modifies the soft label. Higher $\alpha$ values intensify the rectification applied to the prediction probabilities of CLIP. Figure 5(c) shows the average performance of CoOp with ATLaS-ISL across 11 datasets for $\theta$ values in {0.1, 0.2,…, 1}. As $\alpha$ increases, the base-class accuracy gradually improves, while the new-class accuracy initially increases, then declines gently. The harmonic mean accuracy remains generally stable across this range.
Impact of the alternating period $K$. The alternating period $K$ regulates the ratio between epochs using one-hot labels and soft labels for prompt tuning. Figure 5(d) shows the average performance of CoOp with ATLaS-ISL across 11 datasets for $K$ in {1, 2,…, 5}. Note that when $K=1$, the alternating training mechanism is disabled, and only instance-wise soft labels are used for training. As $K$ increases, the base-class accuracy gradually improves, while the new-class accuracy decreases, which aligns with our expectation since a larger $K$ will increase the proportion of one-hot labels used during training. Notably, the harmonic mean accuracy with the alternating mechanism (i.e., $K>1$) consistently outperforms that without the alternating mechanism (i.e., $K=1$), validating the effectiveness of the proposed alternating mechanism.
D.2 Analysis of the Soft Labels Generated by CSL and ISL
Here we provide a case study with detailed explanations in Figure A6, showing the soft labels generated by CSL and ISL for two same-class images. Firstly, CSL enhances vanilla LS by incorporating inter-class similarity knowledge. While vanilla LS assigns equal value to non-target classes, CSL generates more informed soft labels (e.g., 0.96 confidence for target class, with varying confidence for non-target classes based on their similarities). Secondly, ISL generates instance-specific soft labels to capture intra-class variance (e.g., 0.93 and 0.60 confidence for the target class of different images). For an image containing both a human and a beagle, ISL effectively captures this contextual information and assigns 0.24 confidence to the ‘boxer’ class, showing its ability to generate more semantically rich soft labels.
D.3 Combination of CSL and ISL
We explored CSL and ISL mixing ($w\times\mathrm{CSL}+\left(1-w\right)\times\mathrm{ISL}$). As shown in Figure A7, on ImageNet, ATLaS achieves the best performance at $w=0.3$, surpassing individual CSL and ISL. However, we found that this mixing approach does not outperform standalone ISL or CSL across all scenarios.
Figure A6: Case study of the soft labels generated by ISL and CSL.
Figure A7: ATLaS with mixed ISL and CSL on the ImageNet dataset under the base-to-new generalization setting.
D.4 Futher Analysis of the Alternating Period $K$
As mentioned in Sec. 3.3, vanilla LS uses a uniform distribution to generate low-quality soft labels ($\kappa>1$). In such case, the alternating mechanism can improve the performance of vanilla LS. Figure A8 shows that when the alternating training mechanism is used (i.e., $K>1$), ATLaS maintains advantages over vanilla LS.
However, when high-quality soft labels (e.g., ISL) are used, $\kappa$ can be less than 1. Under this condition, Theorem 1 indicates that non-alternating LS achieves a tighter convergence bound than ATLaS. Therefore, Figure 5(d) also aligns with our theoretical analysis, where $K=1$ (using ISL without alternating training) achieves better new-class accuracy compared to other $K$ using the alternating mechanism. Nevertheless, the alternating mechanism can enhance the overall performance. Figure 5(d) shows that the highest H score, which reflects balanced effectiveness on both Base and New classes, is achieved at $K=2$. This occurs because ATLaS-ISL improves base-class accuracy through one-hot label training while preventing overfitting to base classes.
Figure A8: The average performance of ATLaS on all datasets with different $K$ under the base-to-new generalization setting
D.5 Computational Cost
Table A10: Computational cost of the proposed ATLaS, where the model is trained on the base classes of ImageNet with 16-shot examples per class.
Method Train. time Infer.time H (Avg.)
CoOp 1746s 2.33ms 71.66
+ATLaS +0s +0ms 72.30
+ATLaS-CSL +1.96s +0ms 74.39
+ATLaS-ISL +2.47s +0ms 76.79
[…]
Appendix E Limitations
While ATLaS demonstrates significant improvements in enhancing the generalization performance of prompt tuning for vision-language models, it is worth noting that the effectiveness of ATLaS and its variants (i.e., ATLaS-CSL and ATLaS-ISL) may vary across different settings (e.g., cross-dataset transfer, domain generalization) and integration with different prompt tuning methods (e.g., CoOp, MaPLe). This variability suggests that practitioners might need to empirically evaluate all variants to identify the optimal choice for a new setting or architecture. We will design more universally effective soft labels in future work.

### v2

@@ V Conclusion
Acknowledgments
This work is supported by NSFC grant (under No. 62136005).
    Appendix I — Proof of Lemma
    • I-A — Proofs
    • I-B — Technical Lemma 2
    • I-C — Proof of Theorem 1
    • I-D — Convergence Analysis for Label Smoothing
@@ Acknowledgments
    Appendix II — Implementation Details
@@
    Appendix III — Experimental Details
    • III-A — Full Results of the Base-to-New Generalization Setting
I Proofs
@@
I-A Proof of Lemma 1
Recall Lemma 1 in Section III:
Lemma 1
@@ I-B Technical Lemma 2
I-C Proof of Theorem 1
Here we first restate Theorem 1 in Section III as follows:
Theorem 1
@@ I-C Proof of Theorem 1
Proof. Recall Assumption 1 says the objective function $F(\mathbf{v})$ is $\beta$-smooth, i.e., $\left|\nabla F(\mathbf{v})-\nabla F(\mathbf{u})\left|\leq\beta\right|\mathbf{v}-\mathbf{u}\right|$, which has an equivalent expression [58]:
@@
I-D Convergence Analysis for Label Smoothing
In this section, we include the convergence analysis for label smoothing derived by [39]. Note here the context vector $\mathbf{v}_{t}$ are the only learnable parameters, slightly different from the version in [39] where all the parameters $\mathbf{x}_{t}$ in the model are learnable.
Lemma 3
@@ II Implementation Details
The textual prompts for different datasets used in CSL and ISL follow [1, 16], as shown below:
@@
III Experimental Details
Due to the page limit in the main body, we provide detailed results in this section.
III-A Full Results of the Base-to-New Generalization Setting
In this section, we provide the base-to-new generalization performance on 11 datasets and additionally integrate our ATLaS and its variants with the state-of-the-art PromptSRC method [31]. The results in Tables A1, A2 and A3 show that the proposed ATLaS method can consistently improve the new-class accuracy and harmonic mean accuracy over baseline methods, highlighting its effectiveness in enhancing prompt generalization. However, we observe limited performance gains when applying ATLaS to PromptSRC, potentially attributable to both methods leveraging CLIP’s multi-modal capabilities to capture and exploit inter-class and instance-class relationships. Nonetheless, the proposed ATLaS method achieves performance improvements for PromptSRC on all the datasets.